OpenStudy (anonymous):

Verify the identity. cos (x - y) - cos (x + y) = 2 sin x sin y

OpenStudy (anonymous):

is this a recipricol identity??

OpenStudy (anonymous):

cos (x - y) = cosx cosy +sinxsiny cos (x +y) = cosx cosy -sinxsiny

OpenStudy (anonymous):

thus cos (x - y) - cos (x + y) = (cosx cosy +sinxsiny ) - (cosx cosy -sinxsiny ) =cosx cosy +sinxsiny -cosx cosy +sinxsiny =2sinxsiny